OpenStudy (fools101):

Find the simpilest polynomial function with the given zeros. -4, -1, 2 A) P(x) = x3 + 3x2 - 6x - 8 B) P(x) = 2x3 - 3x2 - 6x - 2 C) P(x) = 4x3 + 2x2 - 6x + 8 D) none of the above

OpenStudy (fools101):

please help!

OpenStudy (kira_yamato):

P(x) = (x + 4)(x - 2)(x + 1) = (x - 2)(x^2 + 5x + 4) = x^3 + 3x^2 - 6x - 8 (A)

OpenStudy (anonymous):

(x+4)(x+1)(x-2) = x^3 +3x^2-6x-8 A

OpenStudy (fools101):

THANK U!!!! KIRA AND JIMMYREP!!!! : )